# CITADEL: A Semi-Supervised Active Learning Framework for Malware Detection Under Continuous Distribution Drift

### v1

@@ VII Discussion
Our findings demonstrate that semi-supervised learning with active sample selection is a viable alternative to fully supervised methods for malware detection in drift-heavy environments. Unlike existing SOTA [11, 20] methods that rely on semantic structures (e.g., API graphs) within malware, our approach is designed to remain effective even in the absence of such structures, particularly under long-term evaluation scenarios. Further, CITADEL consistently performs well even with limited labels and achieves high computational efficiency. However, performance gains taper off beyond a 40% label ratio, and classification remains challenging for singleton and ambiguous samples. Contrastive loss improves boundary separation, but additional techniques like denoising or adaptive representation learning may be beneficial.
VIII Conclusion
@@
In this work, we introduce CITADEL, a semi-supervised active learning framework for Android malware detection under long-term concept drift. CITADEL combines malware domain focused augmentation strategies tailored to binary feature vectors with a multi-criteria active sample selection mechanism. Through extensive evaluations on four real-world Android malware datasets, we demonstrate that CITADEL consistently outperforms state-of-the-art baselines. Notably, on the most temporally drifted LAMDA dataset, CITADEL achieves up to a 14% improvement in F1 score while requiring only 40% of the labeled data. It also delivers substantial efficiency gains, with up to faster training and fewer operations. We believe that semi-supervised and active learning-based approaches like CITADEL hold strong promise for real-world deployment due to their robustness, efficiency, and adaptability in the face of evolving malware threats.
@@ Appendix A Family and Singleton Statistics in the Datasets
| APIGraph [13] | 104 | 36 | 172 | 68 | 175 | 55 | 193 | 53 | 199 | 68 | 147 | 48 | 128 | 45 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| Chen-AZ [11] | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 121 | 42 | 82 | 31 | 51 | 17 | - | - | - | - | - | - | - | - |
| LAMDA [15] | - | - | 213 | 1320 | 231 | 2046 | - | - | 375 | 4819 | 207 | 7324 | 373 | 16531 | 635 | 15396 | 588 | 24791 | 295 | 24083 | 651 | 20147 | 224 | 4839 | 64 | 538 | 8 | 18 |
| MaMaDroid [14] | - | - | 1 | 32 | 1 | 47 | - | - | 96 | 132 | 48 | 391 | 80 | 521 | 134 | 291 | 117 | 455 | 62 | 805 | 130 | 515 | 93 | 717 | 57 | 526 | 7 | 11 |
@@
Table VII presents the year-wise counts of singleton malware samples—those appearing only once—for the same datasets. APIGraph reports singletons from 2012–2018, peaking at 68 in 2013, with no data afterward. Chen-AZ includes singleton counts from 2019–2021, peaking at 42 in 2019. LAMDA shows extensive singleton data from 2013–2025, with a major surge between 2018 and 2022, peaking at 24,791 in 2020. Dashes indicate years without data. These singleton trends reflect temporal variations and dataset-specific dynamics, potentially signaling malware evolution or labeling noise. Additionally, Table VII also summarizes the year-wise counts of malware families across the four datasets. APIGraph lists family totals from 2012–2018, increasing from 104 families in 2012 to a peak of 199 in 2016, followed by a gradual decline. Chen-AZ provides family counts only for 2019–2021, ranging from 51 to 121 during this period. LAMDA’s family counts span from 2013 to 2025, starting with 213 families in 2013 and increasing over the years, reaching as high as 651 families in 2022. MaMaDroid reports family data for the same years as LAMDA, but its family count is relatively less. It reached a maximum of 134 families in a single year, and its lowest is one family in 2013–2014. Among all datasets, LAMDA exhibits the broadest temporal coverage and the highest family diversity.
Appendix B CITADEL vs Chen-AL performance visualization
Figures 6, 7, 8, and 9 visualize monthly trends in F1-score, false positive rate (FPR), and false negative rate (FNR) for CITADEL versus Chen-AL [11] on the APIGraph, Chen-AZ (AndroZoo), LAMDA, and MaMaDroid datasets, respectively.
@@ Appendix C CITADEL Hyperparameter tuning
To analyze the impact of augmentation probability on performance, we conducted a systematic tuning experiment over both weak and strong augmentations in CITADEL. Specially, we employ random bit flip from Bernoulli distribution with probability to determine the impact of feature perturbation and fine-tune to get optimal values. We evaluate three settings: (i) fixing the strong augmentation probability () while varying the weak augmentation ; (ii) fixing the weak augmentation probability () while varying the strong augmentation ; and (iii) applying the same augmentation probability for both weak and strong augmentation. As shown in Table IX, the performance remains stable when is small (e.g., 1–5%), but gradually degrades as increases beyond this range.
| Applied aug () | 0.01 | 0.05 | 0.1 | 0.2 | 0.3 | 0.4 | 0.5 | 0.6 | 0.7 |
| Strong Fixed () | 93.5 | 93.5 | 93.4 | 93.3 | 93.1 | 93.1 | 92.8 | 92.1 | 91.7 |
| Vary Weak () | 92.8 | 93.5 | 93.4 | 93.3 | 93.2 | 93.3 | 93.1 | 92.6 | 92.5 |
| Same Weak and Strong () | 93.2 | 93.5 | 93.4 | 93.3 | 93.2 | 93.3 | 93.1 | 92.4 | 91.8 |
@@ Appendix D CITADEL performance under label noise
Experiments. To evaluate the robustness of CITADEL against noisy supervision, we conduct a controlled experiment using the APIGraph dataset. We start with cleanly labeled Android malware samples and progressively inject synthetic label noise by randomly flipping a portion of the ground-truth labels to incorrect ones. The label noise rate varies from 0% to 90% in increments of 10%. The training data may contain misclassified or mislabeled samples, which is common in large-scale malware datasets due to incorrect labels provided by automated anti-virus engines.
Results. CITADEL demonstrates a strong tolerance to label corruption. As shown in Table X, CITADEL maintains an F1-score above 88% for up to 20% label noise, and even with extreme corruption levels of 90% noisy labels, it achieves an F1-score of approximately 70%. This corresponds to only a 23% relative drop compared to the noise-free case. These results highlight that CITADEL retains stable, even under heavily corrupted supervision.
@@ Appendix E CITADEL with curriculum learning
Experiments. To evaluate whether Curriculum Learning (CL) can mitigate malware drift, we designed a baseline experiment where training samples were ordered from easy to hard following the CL principle proposed by [44]. Unlike Active Learning (AL), which adaptively selects new informative samples during training, CL assumes a smooth progression in task difficulty. This assumption rarely holds in malware detection, where new variants appear abruptly. We used the 2012 training subset of the APIGraph dataset and ranked all instances by instance hardness following [44]. To simulate limited human annotation effort, we additionally labeled 400 hard samples per month as in [59]. Both settings used identical hyperparameters and data splits as our main experiments for fair comparison with active learning.
@@
Results. Table XI summarizes the comparison. Without CL, the model achieved an F1 score of 71.2%, with relatively high FNR and FPR. Incorporating CL and limited hard-sample feedback improved the F1 score to 83.0%, showing moderate gains in generalization and convergence stability. However, our proposed Active Learning (CITADEL) approach—jointly considering margin, distance, and confidence (Table IV)—achieved a substantially higher 93.5% F1, with a much lower FNR of 8.4% and FPR of 0.4%. These results demonstrate that while CL helps on static datasets, AL is far more effective under evolving malware conditions due to its adaptability and targeted sample selection.
@@ Appendix F Experiments on EMBER Feature
Experiments. On the EMBER dataset, we use samples from 2018-01 to 2018-02 for initial training, and begin active learning from 2018-03 to 2018-12 using different methods—Chen-AL [11], CADE [20], and CITADEL. For this feature space, since Bernoulli bit-flip is not applicable, we employ Bernoulli feature masking as the augmentation strategy in CITADEL. This approach regularizes the model by randomly masking input features, thereby improving generalization under limited labeling budgets.
Results. Table XII presents the performance of CADE, Chen-AL, and CITADEL under different labeling budgets on the EMBER dataset. CITADEL consistently achieves the best results across all budgets. At a budget of 50, CITADEL attains an F1-score of 87.9%, outperforming CADE (85.4%) and Chen-AL (83.4%). With a budget of 400, CITADEL further improves to an F1-score of 89.4%, while CADE reaches 86.0%.

### v2

@@ 7 Discussion
Our results suggest that semi-supervised active learning is a practical alternative to fully supervised malware detection under long-term distribution drift. Unlike prior methods that rely on semantic structures [chen2023continuous, yang2021cade], CITADEL remains effective even even such structures are absent, particularly under long-term evaluation scenarios. CITADEL consistently performs well with limited labeled data and achieves high computational efficiency. However, we observe diminishing gains beyond a 40% label ratio, and classification remains challenging for singleton and ambiguous samples.
Recommendation for Practitioners.
Based on our experience in developing CITADEL, we outline practical recommendations for adapting CITADEL. First, models should be initialized with reliable labeled data, as accurate early supervision supports stable downstream adaptation. Second, our results indicate that mild augmentation improves performance, whereas excessive augmentation can degrade performance (see Appendix C.1 for sensitivity analysis). We further recommend using a high pseudo-label confidence threshold (), since reducing this threshold is found to negatively impact performance.
Third, class imbalance should be monitored carefully throughout training and retraining to avoid bias toward the majority class. Fourth, at scale, efficient training benefits from standard high-performance setups; including batch-wise computation with PyTorch tensors on CUDA-enabled GPUs; our replication artifacts follow this design. Finally, we recommend using CITADEL’s Multi-Criteria sample selection strategy since it prioritizes informative and ambiguous samples for labeling. In practice, integrating this selector with periodic human review supports continual adaptation to evolving drift while maintaining strong detection performance.
ML Attacks against CITADEL.
As with other adaptive ML-based malware detectors, CITADEL operates in adversarial settings where evasion and poisoning attacks are well known challenges. In particular, the use of pseudo-labeling and selective querying can, in principle, be targeted by clean-label poisoning attempts that aim to gradually influence the decision boundaries. Our multi-criteria sample selection is designed to mitigate such risks by prioritizing ambiguous and informative samples rather than relying on a single heuristic. However, further strengthening against adversarial manipulation remains an active direction of future work.
8 Conclusion
@@
In this work, we present CITADEL, a semi-supervised active learning framework for Android malware detection under long-term concept drift. CITADEL combines domain-specific feature augmentations for binary malware features vectors with a multi-criteria active sample selection mechanism. Extensive evaluations on four real-world Android malware datasets show that CITADEL consistently outperforms state-of-the-art baselines. On the most temporally drifted LAMDA dataset, CITADEL achieves up to a 14% F2 improvement while using only 40% of the labeled data. It also provides substantial efficiency gains, with up to faster training and fewer operations. These results suggest that semi-supervised active learning-based approaches such as CITADEL hold strong potential for real-world deployment due to their robustness, efficiency, and adaptability in the face of evolving malware threats.
@@ Appendix A Family and Singleton Statistics in the Datasets
| APIGraph [api_graph_dataset] | 104 | 36 | 172 | 68 | 175 | 55 | 193 | 53 | 199 | 68 | 147 | 48 | 128 | 45 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| Chen-AZ [chen2023continuous] | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 121 | 42 | 82 | 31 | 51 | 17 | - | - | - | - | - | - | - | - |
| LAMDA [haque2025lamda] | - | - | 213 | 1320 | 231 | 2046 | - | - | 375 | 4819 | 207 | 7324 | 373 | 16531 | 635 | 15396 | 588 | 24791 | 295 | 24083 | 651 | 20147 | 224 | 4839 | 64 | 538 | 8 | 18 |
| MaMaDroid [mamadroid] | - | - | 1 | 32 | 1 | 47 | - | - | 96 | 132 | 48 | 391 | 80 | 521 | 134 | 291 | 117 | 455 | 62 | 805 | 130 | 515 | 93 | 717 | 57 | 526 | 7 | 11 |
@@
Table 7 presents the year-wise counts of singleton malware samples—those appearing only once—for the same datasets. APIGraph reports singletons from 2012–2018, peaking at 68 in 2013, with no data afterward. Chen-AZ includes singleton counts from 2019–2021, peaking at 42 in 2019. LAMDA shows extensive singleton data from 2013–2025, with a major surge between 2018 and 2022, peaking at 24,791 in 2020. Dashes indicate years without data. These singleton trends reflect temporal variations and dataset-specific dynamics, potentially signaling malware evolution or labeling noise. Additionally, Table 7 also summarizes the year-wise counts of malware families across the four datasets. APIGraph lists family totals from 2012–2018, increasing from 104 families in 2012 to a peak of 199 in 2016, followed by a gradual decline. Chen-AZ provides family counts only for 2019–2021, ranging from 51 to 121 during this period. LAMDA’s family counts span from 2013 to 2025, starting with 213 families in 2013 and increasing over the years, reaching as high as 651 families in 2022. MaMaDroid reports family data for the same years as LAMDA, but its family count is relatively less. It reached a maximum of 134 families in a single year, and its lowest is one family in 2013–2014. Among all datasets, LAMDA exhibits the broadest temporal coverage and the highest family diversity.
Appendix B Visual Analysis of CITADEL vs Chen-AL Performance
Figures 6 visualize monthly trends in F1-score, false positive rate (FPR), and false negative rate (FNR) for CITADEL versus Chen-AL [chen2023continuous] on the APIGraph, Chen-AZ (AndroZoo), LAMDA, and MaMaDroid datasets, respectively. In addition, we provide an in-depth comparison between Chen-AL [chen2023continuous] and CITADEL in terms of the total count of False Negative (FN) and False Positive (FP) over the year on LAMDA dataset [haque2025lamda] in Table 8 and Figure 7. Clearly, after 2017, Chen-AL [chen2023continuous] fails to generalize between malware and benign samples. The low false-negative (FN) count and high false-positive (FP) count relative to the total number of samples indicate that the model increasingly classifies benign samples as malware, resulting in an effectively zero FN rate. Consequently, Chen-AL does not cope with dynamic decision boundary shifts induced by longitudinal concept drift, particularly on the LAMDA dataset [haque2025lamda]. Furthermore, substantial label drift causes a significant performance degradation in prior techniques, whereas CITADEL consistently retains its performance even under these challenging conditions.
| (a) APIGraph | ||
| (b) Chen-AZ | ||
| (c) LAMDA | ||
| (d) MaMaDroid | ||
| Year | Total | Chen-AL [chen2023continuous] | CITADEL | ||
| FN | FP | FN | FP | ||
| 2014 | 8652 | 586 | 167 | 439 | 178 |
| 2016 | 21839 | 1324 | 8646 | 1796 | 1458 |
| 2017 | 19829 | 0 | 15557 | 525 | 354 |
| 2018 | 20859 | 0 | 12989 | 1123 | 479 |
| 2019 | 18210 | 0 | 9893 | 1573 | 225 |
| 2020 | 20415 | 0 | 11144 | 967 | 161 |
| 2021 | 16231 | 0 | 9106 | 496 | 101 |
| 2022 | 17284 | 0 | 8954 | 856 | 89 |
| 2023 | 10871 | 0 | 9293 | 302 | 46 |
| 2024 | 9686 | 0 | 9527 | 57 | 42 |
| 2025 | 8933 | 0 | 8928 | 4 | 34 |
@@ Appendix C CITADEL Hyperparameter tuning
C.1 Augmentation Sensitivity Analysis
Experimental Setup.
We investigate the impact of weak () and strong () perturbation probabilities in CITADEL’s Bernoulli distribution–based augmentation on the APIGraph dataset under an active learning setting with a fixed labeling budget of 400 samples. We first conduct a coarse sweep in which the weak and strong perturbation probabilities are set to the same value () and varied from 0.0 to 0.9 at intervals of 0.1 to capture overall performance trends. Based on the observed behavior, we then perform a fine-grained sweep over a narrower range () to more precisely identify favorable augmentation strengths.
To further examine the individual effects of weak and strong augmentations, we next control their probabilities independently. We fix the strong perturbation probability at a high value () while varying the weak perturbation probability from 0.00 to 0.09. We then repeat this experiment with a low strong perturbation probability (), again varying over the same range. The results are reported in Table 10 and illustrated in Figure 8.
Analysis.
Across all configurations, introducing small amounts of perturbation consistently improves performance, with F1 scores increasing from 90.2% to a maximum of 93.5%. As the weak perturbation probability increases further, performance gradually declines. This degradation occurs because excessive weak augmentation introduces substantial noise into both labeled and unlabeled samples, which can distort class semantics and destabilize learning. In contrast, mild weak perturbations improve generalization by promoting robustness while preserving discriminative structure.
Strong augmentation exhibits a different trend. Even at higher strong perturbation probabilities, performance degradation is less severe. This behavior arises because strong augmentation is applied exclusively to the unsupervised branch, and CITADEL employs a confidence-based filtering mechanism that prevents heavily distorted samples from contributing to the consistency objective. Overall, these results highlight the importance of separately controlling weak and strong perturbations. The best performance is achieved with a small weak perturbation probability and a low strong perturbation probability—specifically, and —which yields an effective balance between robustness and discriminative learning.
| Same Weak and Strong Perturbation () | ||||||||||
| 0.00 | 0.10 | 0.20 | 0.30 | 0.40 | 0.50 | 0.60 | 0.70 | 0.80 | 0.90 | |
| F1 | 90.230.31 | 91.740.27 | 86.960.39 | 80.940.48 | 77.920.52 | 14.620.22 | 11.080.19 | 9.030.45 | 7.880.39 | 6.590.32 |
| 0.00 | 0.01 | 0.02 | 0.03 | 0.04 | 0.05 | 0.06 | 0.07 | 0.08 | 0.09 | |
| F1 | 90.230.31 | 93.490.20 | 93.450.21 | 93.380.23 | 93.340.24 | 93.300.25 | 93.210.22 | 93.140.27 | 92.890.29 | 92.650.31 |
| Varying Weak, Strong Fixed | ||||||||||
| 0.00 | 0.01 | 0.02 | 0.03 | 0.04 | 0.05 | 0.06 | 0.07 | 0.08 | 0.09 | |
| F1 () | 90.630.28 | 92.800.24 | 92.740.25 | 92.650.26 | 92.530.27 | 92.160.29 | 92.060.30 | 91.840.31 | 91.240.32 | 91.160.33 |
| F1 () | 90.330.29 | 93.540.23 | 93.420.24 | 93.290.25 | 93.180.26 | 93.180.27 | 93.020.28 | 92.560.30 | 92.360.31 | 92.300.32 |
C.2 Threshold and Coefficients
@@ Appendix D CITADEL Performance under Label Noise
Experiments.
To evaluate the performance of CITADEL against noisy supervision, we conduct a controlled experiment using the APIGraph dataset. We start with cleanly labeled Android malware samples and progressively inject synthetic label noise by randomly flipping a portion of the ground-truth labels to incorrect ones. The label noise rate varies from 0% to 90% in increments of 10%. The training data may contain misclassified or mislabeled samples, which is common in large-scale malware datasets due to incorrect labels provided by automated anti-virus engines.
Results.
CITADEL demonstrates a strong tolerance to label corruption. As shown in Table 11, CITADEL maintains an F1-score above 88% for up to 20% label noise, and even with extreme corruption levels of 90% noisy labels, it achieves an F1-score of approximately 70%. This corresponds to only a 23% relative drop compared to the noise-free case. These results highlight that CITADEL retains stable, even under heavily corrupted supervision.
@@ Appendix E CITADEL with Curriculum Learning
Experiments.
To evaluate whether Curriculum Learning (CL) can mitigate malware drift, we designed a baseline experiment where training samples were ordered from easy to hard following the CL principle proposed by [zhou2020curriculum]. Unlike Active Learning (AL), which adaptively selects new informative samples during training, CL assumes a smooth progression in task difficulty. This assumption rarely holds in malware detection, where new variants appear abruptly. We used the 2012 training subset of the APIGraph dataset and ranked all instances by instance hardness following [zhou2020curriculum]. To simulate limited human annotation effort, we additionally labeled 400 hard samples per month as in [kong2021adaptive]. Both settings used identical hyperparameters and data splits as our main experiments for fair comparison with active learning.
Results.
Table 12 summarizes the comparison. Without CL, the model achieved an F1 score of 71.2%, with relatively high FNR and FPR. Incorporating CL and limited hard-sample feedback improved the F1 score to 83.0%, showing moderate gains in generalization and convergence stability. However, our proposed Active Learning (CITADEL) approach—jointly considering margin, distance, and confidence (Table IV)—achieved a substantially higher 93.5% F1, with a much lower FNR of 8.4% and FPR of 0.4%. These results demonstrate that while CL helps on static datasets, AL is far more effective under evolving malware conditions due to its adaptability and targeted sample selection.
@@ Appendix F Experiments on EMBER Feature
Experiments.
On the EMBER dataset [anderson2018ember], we use samples from 2018-01 to 2018-02 for initial training, and begin active learning from 2018-03 to 2018-12 using different methods—Chen-AL [chen2023continuous], CADE [yang2021cade], and CITADEL. For this feature space, since Bernoulli bit-flip is not applicable, we employ Bernoulli feature masking as the augmentation strategy in CITADEL. This approach regularizes the model by randomly masking input features, thereby improving generalization under limited labeling budgets.
Results.
Table 13 presents the performance of CADE, Chen-AL, and CITADEL under different labeling budgets on the EMBER dataset. CITADEL consistently achieves the best results across all budgets. At a budget of 50, CITADEL attains an F1-score of 87.9%, outperforming CADE (85.4%) and Chen-AL (83.4%). For budget 400, CITADEL further improves to an F1-score of 89.4%, while CADE reaches 86.0%.
@@ Results.
Appendix G Computational Resources
All experiments are conducted on a server with dual-socket Intel Xeon Gold 6430 CPUs, providing 64 physical cores (128 logical cores) and 1TB of main memory. The server is equipped with 4 NVIDIA H100 NVL GPUs, each with 94GB of GPU memory. The software environment uses CUDA 12.8 with NVIDIA driver version 570.124.X. Each run of CITADEL uses approximately 3–4GB of main memory and 2–3 GB of GPU memory. CITADEL can also be executed on cloud-based platforms such as Google Colab, where an NVIDIA A100 or H100 GPU is recommended for acceleration.